OpenStudy (anonymous):

EXpand the following expression so that there is no subtraction in the domain of the sine function: sin(2x-5) Show all work please.

OpenStudy (dumbcow):

\[\sin(a-b) = \sin a \cos b - \cos a \sin b\]

OpenStudy (anonymous):

I got sin(2x)cos(5)-sin(5)cos(2x)

OpenStudy (dumbcow):

yes